OpenStudy (anonymous):

If a spherical baseball has a diameter of 7 cm, calculate the volume of the ball.

OpenStudy (anonymous):

\[V=4/3*\pi*x^3\] \[V=1438.8 cm^3\]

OpenStudy (anonymous):

The volume of a sphere is: \[\frac{4}{3}\pi r^3\] in this case r is 3.5 cm so the volume of the baseball is: \[\frac{4}{3}\pi (3.5)^3 = 179.6 cm^3\] to 4 sig figs.

OpenStudy (anonymous):

I think the solution by PJW may have been calculated with r = 7.